OpenStudy (anonymous):

mike wants to make a meatloaf. His recipe uses a total of 7 pounds of meat. If he uses a 3 to 1 ratio of beef to pork, how much pork will he use ? Enter your answer as a mixed number in simplest terms.

OpenStudy (dumbcow):

ahh i see. so he uses 3 times more beef than pork B = 3P B + P = 7 --> 3P +P = 7 --> 4P = 7 --> P = 7/4 B = 3*7/4 = 21/4